# On the solvable-unsolvable transition due to noise-induced chaos in digital memcomputing

### v1

@@ 4 Discussion
In this study, we employed various indicators to examine how numerical and physical noise impact the behavior of digital memcomputing machines. From the practical perspective, the influence of both types of noise appears similar: with increasing noise strength, the machines’ problem-solving capability diminishes from 100% to a lower percentage, eventually to 0%. This is evident in Fig. 1 for numerical noise and Fig. 4 for physical noise.
@@
To analyze the loss of solvability, we employed the Lyapunov exponent analysis. We utilized the algorithm by Benettin et al. [31] in the study of numerical noise, and a conventional algorithm to handle physical noise, leading to the determination of the largest Lyapunov exponents [^2] We note that the Lyapunov exponent analysis was focused exclusively on the fast degrees of freedom (the logical variables). We believe that these are adequate for assessing whether the overall system dynamics (logical variables combined with memory variables) exhibit chaotic behavior.. The averaged LLEs are presented in Fig. 1 for numerical noise and in Fig. 4 for physical noise. Both scenarios exhibit a steady increase in the averaged LLEs; however, with numerical noise, this increase slightly precedes the decline in the solved instance percentage, whereas with physical noise a delay is observed. The reason for this observation is a possible underestimation of LLE by the standard algorithm that evaluates a single exponent versus the spectra in the algorithm by Benettin et al. [31]. Additionally, the standard algorithm was applied to a modified system (without constraints on the $v_{n}$’s) that may have a slightly different behavior. Based on the LE analysis, we have concluded that the solvable-unsolvable transition in digital memcomputing machines can be associated with a noise-induced chaos.
@@
The power spectra found in this work (see, e.g., Figs. 2-4) provide additional insight into the dynamics of digital memcomputing machines. In the case of numerical noise, we have noticed that all instances go through four stages when increasing the noise level: at small time steps, they exhibit either no chaos or weak “transient chaos” by showing very little or no chaotic behavior; as the time step increases, they enter the stage of “transient chaos” while still being able to reach the solution; and by increasing the time step further the system becomes fully chaotic and does not solve anymore. In the final stage, when $\Delta t$ is very large, the power spectra include spikes indicating a strong periodic component. It was observed that at very large time steps, the logical variables oscillate between ±$\pm 1$. In the case of physical noise, all stages but the last one are also present. This difference may be explained by the fact that unlike physical noise, numerical noise is “non-local” in time.
5 Conclusions
@@
In conclusion, our investigation into the noisy dynamics of digital memcomputing machines (DMMs) demonstrates a fairly high robustness when solving problems in the presence of numerical or physical noise. The effects of noise have been observed through several indicators, such as the percentage of solved instances, power spectra, and Lyapunov exponents. We have observed that the effects of numerical and physical noise are overall qualitatively similar. The largest difference has been found in the case of very strong noise, which, in the numerical case, leads to quasi-periodic-like behavior (in the form of well distinguishable spikes in the power spectra), while strong physical noise results in broadband spectra. Additionally, we emphasize that there exists a “transiently chaotic” regime, characterized by positive mean largest Lyapunov exponents, where DMMs still demonstrate solvability. These results shed further light on this new class of computing machines and could be used to control whether they operate in the optimal dynamical regime.
Acknowledgement
@@
Appendix A Time evolution of largest Lyapunov exponents
In Figure A1, we plot the LLE as a function of time throughout the dynamics for cases with both little and large numerical noise. The trajectories exhibit large spikes in LLE, which we attribute to the existence of saddle points in the phase space. The unstable directions of the saddle points would naturally produce positive LLEs.
(a) Refer to caption(b) Refer to caption
@@ Appendix A Time evolution of largest Lyapunov exponents
Appendix B $1/f$ noise
As mentioned in the main text, in the case of physical noise there is an additional feature present in the power spectra at very low frequencies (Inset (2) in Fig. 4). We attribute this feature to flicker noise [41]. This phenomenon is characterized by a power spectral density that scales inversely with frequency, following the form $1/f^{\alpha}$, with typically $1\leq\alpha\leq 2$.
(a) Refer to caption (b) Refer to caption
Figure A2: Log-log plot of the power spectrum for the time derivative of the $v_{n}$’s. We simulated a problem instance with $N=1000$, with ratio $M/N=7$. In any stochastic process, there is a $1/f$ power spectral density down to the lowest frequency $f$ whenever $f>\Omega_{min}$, with $\Omega_{min}$ being the minimum relaxation frequency [42]. Since the minimum relaxation frequency of a system is generally fixed, the power spectrum deviates from the $1/f$ trend for frequencies $f$ smaller than $\Omega_{min}$, marking a crossover to a different spectral regime. On (a), the power spectrum is averaged over 100 realizations with noise strength $\Gamma=1$, while for (b) the noise strength is $\Gamma=2$ and the same average is performed.
This is illustrated in the power spectra of Fig. 4. The characteristic $1/f$ behavior is confirmed by the spectral slopes shown in Fig. A2, observed for both intermediate and high noise strength. Indeed, at low frequencies, the spectrum scales as $f^{-1.23}$ for $\Gamma=1$ and as $f^{-1.46}$ for $\Gamma=2$. The noise saturates at frequencies comparable to the lowest relaxation frequency of the process [42].
@@ Appendix B $1/f$ noise
[…]
[1]organization=Department of Physics and Astronomy, University of South Carolina, city=Columbia, SC, postcode=29208, country=USA
[…]
[2]organization=École Polytechnique Fédérale de Lausanne, city=Lausanne, postcode=1015, country=Switzerland
[…]
[3]organization=Department of Physics, University of California, San Diego, city=La Jolla, CA, postcode=92093-0319, country=USA

### v2

@@ IV Discussion
In this study, we employed various indicators to examine how numerical and physical noise impact the behavior of digital memcomputing machines. From a practical perspective, the influence of both types of noise appears similar: with increasing noise strength, the machines’ problem-solving capability diminishes from 100% to a lower percentage, eventually to 0%. This is evident in Fig. 1 for numerical noise and Fig. 4 for physical noise.
@@
To analyze the loss of solvability, we employed the Lyapunov exponent analysis. We utilized the algorithm by Benettin et al. Benettin et al. (1980) in the study of numerical noise, and a conventional algorithm to handle physical noise, leading to the determination of the largest Lyapunov exponents [^2] We note that the Lyapunov exponent analysis was focused exclusively on the fast degrees of freedom (the logical variables). We believe that these are adequate for assessing whether the overall system dynamics (logical variables combined with memory variables) exhibit chaotic behavior.. The EAMLLEs are presented in Fig. 1 for numerical noise and in Fig. 4 for physical noise. With numerical noise, a strong increase in the EAMLLE precedes the decline in the solved-instance percentage, whereas with physical noise a slight delay is observed. The reason for this observation is a possible underestimation of LLE by the standard algorithm that evaluates a single exponent versus the spectra in the algorithm by Benettin et al. Benettin et al. (1980). Additionally, the standard algorithm was applied to a modified system (without constraints on the $\textnormal{v}_{n}$’s) that may have a slightly different behavior. Based on the LE analysis, we conclude that the solvable-unsolvable transition in digital memcomputing machines can be associated with a noise-induced chaos.
@@
The power spectra found in this work (see, e.g., Figs. 2-4) provide additional insight into the dynamics of digital memcomputing machines. In the case of numerical noise, we have noticed that all instances go through four stages when increasing the noise level: at small time steps, they exhibit either no chaos or weak “transient chaos” by showing very little or no chaotic behavior; as the time step increases, they enter the stage of “transient chaos” while still being able to reach the solution; and by increasing the time step further the system becomes fully chaotic and does not solve anymore. In the final stage, when $\Delta t$ is very large, the power spectra include spikes indicating a strong periodic component. It was observed that at very large time steps, the logical variables oscillate between $\pm 1$. In the case of physical noise, all stages but the last one are also present. This difference may be explained by the fact that unlike physical noise, numerical noise is “non-local” in time.
V Conclusions
@@
In conclusion, our investigation into the noisy dynamics of digital memcomputing machines (DMMs) demonstrates a fairly high robustness when solving 3-SAT problems in the presence of numerical or physical noise. The effects of noise have been observed through several indicators, such as the percentage of solved instances, power spectra, and Lyapunov exponents. We have observed that the effects of numerical and physical noise are overall qualitatively similar. The largest difference has been found in the case of very strong noise, which, in the numerical case, leads to quasi-periodic-like behavior (in the form of well distinguishable spikes in the power spectra), while strong physical noise results in broadband spectra. Additionally, we emphasize that there exists a “transiently chaotic” regime for both types of noise, characterized by positive mean largest Lyapunov exponents, where DMMs still demonstrate solvability. These results shed further light on this new class of computing machines and could be used to control whether they operate in the optimal dynamical regime.
Supplementary Material
The supplementary material includes additional power spectra that further illustrate the dynamics of memcomputing machines in the presence of numerical and physical noise.
Acknowledgements.
@@
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Visualization of transient chaos in DMM
To illustrate the presence of transient chaos in the dynamics of the DMM, we show the number of integration steps required to reach the solution as a function of the initial value $\textnormal{v}_{0}$ for a single 3-SAT instance in Fig. A1. All other variables were initialized identically and $\textnormal{v}_{0}$ was continuously varied from -1 to 1.
Figure A1: Number of integration steps to solution as a function of initial value of $\textnormal{v}_{0}$ for one instance with $N=1000$ and $M/N=7$ using time step $\Delta t=0.1$. $\textnormal{v}_{0}$ is varied monotonically from 1-1 to 1 with an increment of $10^{-3}$.
Appendix B Time evolution of largest Lyapunov exponents
In Fig.  A2, we plot the LLE as a function of time throughout the dynamics for cases with both small and large numerical noise. The trajectories exhibit large spikes in LLE, which we attribute to the existence of saddle orbits in the phase space. The unstable directions of the saddle orbits would naturally produce positive LLEs.
(a) Refer to caption
(b) Refer to caption
@@ Appendix B Time evolution of largest Lyapunov exponents
Appendix C $1/f$ noise
As mentioned in the main text, in the case of physical noise there is an additional feature present in the power spectra at very low frequencies (Inset (2) in Fig. 4). We attribute this feature to flicker noise Dutta and Horn (2024). This phenomenon is characterized by a power spectral density that scales inversely with frequency, following the form $1/f^{\alpha}$, with typically $1\leq\alpha\leq 2$.
(a) Refer to caption
(b) Refer to caption
Figure A3: Log-log plot of the power spectrum for the time derivative of the $\textnormal{v}_{n}$’s. We simulated a problem instance with $N=1000$, with ratio $M/N=7$. In any stochastic process, there is a $1/f$ power spectral density down to the lowest frequency $f$ whenever $f>\Omega_{min}$, with $\Omega_{min}$ being the minimum relaxation frequency Diaz and Di Ventra (2014). Since the minimum relaxation frequency of a system is generally fixed, the power spectrum deviates from the $1/f$ trend for frequencies $f$ smaller than $\Omega_{min}$, marking a crossover to a different spectral regime. On (a), the power spectrum is averaged over 100 realizations with noise strength $\Gamma=1$, while for (b) the noise strength is $\Gamma=2$ and the same average is performed.
This is illustrated in the power spectra of Fig. 4. The characteristic $1/f$ behavior is confirmed by the spectral slopes shown in Fig. A3, observed for both intermediate and high noise strength. Indeed, at low frequencies, the spectrum scales as $f^{-1.23}$ for $\Gamma=1$ and as $f^{-1.46}$ for $\Gamma=2$. The noise saturates at frequencies comparable to the lowest relaxation frequency of the process Diaz and Di Ventra (2014).
@@ Appendix C $1/f$ noise
Supplementary Information for the manuscript:
On the solvable–unsolvable transition due to noise-induced chaos in digital memcomputing